OpenStudy (anonymous):

How do I solve this: x-x^3=2

OpenStudy (anonymous):

\[x-x^3 = x(1-x^2)\]

OpenStudy (anonymous):

therefore either x=2 or 1-x^2=2

OpenStudy (anonymous):

say \[x \neq, then 1-x^2=2\]

OpenStudy (anonymous):

\[1-x^2=2, will give x^2=-1, which has complex roots \therefore only answer is x=2\]

OpenStudy (anonymous):

Well two doesnt work out. The real answer is -1.52 Is their any way to do this algebraically?

OpenStudy (anonymous):

can you retype the question,please

OpenStudy (anonymous):

Sure, \[x - x ^{3} =2 \]

OpenStudy (anonymous):

just to make sure it is x^3, right?

OpenStudy (anonymous):

Yeah its x cubed

OpenStudy (anonymous):

and are you familiar with complex numbers

OpenStudy (anonymous):

sorry for asking

OpenStudy (anonymous):

We've done last year but im not sure Ill remember. Im not even sure if I have to figure this out alebraically. So Ill just ask a friend tommorow. But thanks alot fo helping.

OpenStudy (anonymous):

i think you take the magnitude of the real and imaginary solution you will have the answer as -1.5, which makes complete sense

OpenStudy (anonymous):

else this equation cant have real roots

OpenStudy (anonymous):

Oh wow that makes sense, I think ive got it from here. Thanks again!!